The weak acid of the buffer system is used to prepare the buffer by reacting with a strong base. The conjugated base is going to be obtained from some of the weak acid in reaction with NaOH 1M. This means you need to calculate the total amount of weak acid needed and the amount of 1.00 M NaOH solution you will have to add to create its conjugate form in solution.   Use 1.0M of each of these acids and basic solutions, H3PO4, HCl, HC2H3O2, NaOH, to create a 100ml buffer solution of pH5.
Answer #1

SOLUTION:

Dear candidate you have given 3 acids and one base to prepare buffers. Out of the 3 acids only HC2H3O2, will form a buffer with NaOH. Other two acids (HCl and H3PO4) are strong acids and will not form buffer with NaOH.

CH3COOH + NaOH --------> CH3COO- + Na+ + H2O

pH = pKa + log[CH3COO-] / [CH3COOH]

[CH3COO-] = amount of NaOH added. Let xM of NaOH is added to CH3COOH. Then:

amount of [CH3COO-] = xM

amount of [CH3COOH] = (1 - x)M

pH = pKa + log[x] / [1 - x]

5 = 4.76 + log[x] / [1 - x] ; Because pKa of CH3COOH is 4.76

0.24 = log[x] / [1 - x]

[x] / [1 - x] = antilog (0.24)

[x] / [1 - x] = 1.74

x = 1.74 ( 1-x)

x = 1.74 - 1.74x

2.74x = 1.74

x = 0.63M

Hence Amount of NaOH required is 0.63M